# UFO [Uncaptioned image]: A UI-Focused Agent for Windows OS Interaction

### v4

@@ 1 Introduction
The advent of Large Language Models (LLMs) Ouyang et al. (2022); Achiam et al. (2023); Touvron et al. (2023) has showcased revolutionary potential in solving complex problems akin to human reasoning, planning, and collaboration Hao et al. (2023); Ding et al. (2023); Talebirad & Nadiri (2023). This development brings humanity closer to the realization of Artificial General Intelligence (AGI) Liu et al. (2023), offering assistance in various tasks in our daily lives and exhibiting a robust and comprehensive level of ability that was previously considered challenging Chen et al. (2023b); Shao et al. (2023); Jin et al. (2023). The advancement towards more sophisticated multi-modal visual LLMs, exemplified by GPT-Vision Zhang et al. (2023); Yang et al. (2023b), introduces a visual dimension to LLMs, expanding their capabilities to encompass a myriad of visual tasks and broadening their scope to address challenges that require visual capability Wu et al. (2023a); Zheng et al. (2024); Wake et al. (2023).
@@
In this context, the Windows operating system (OS) stands out as representative platform for LAMs, due to its high market share in the daily use of computer systems Adekotujo et al. (2020), the presence of versatile applications and GUIs built upon it Ramler et al. (2018), and the complexity of tasks that necessitate long-term planning and interaction across various applications Stallings (2005). The prospect of having a general intelligent agent that can comprehend user requests in natural language, and autonomously interact with the UIs of applications built on Windows is highly appealing. Despite the evident demand for developing VLM agents tailored for Windows OS to fulfill user requests, this direction remains largely unexplored, as most existing agents predominantly focus on smartphones Yang et al. (2023a) or web applications Zheng et al. (2024). This gap in exploration presents a vast, untapped opportunity to develop a Windows-specific agent.
To bridge this gap, we introduce UFO [Uncaptioned image], a specialized UI-Focused Agent designed for seamless interaction with the Windows OS, leveraging the cutting-edge capabilities of the VLM GPT-Vision Yang et al. (2023b). UFO adopts a dual-agent framework, with each agent analyzing screenshots and extracting information from the GUI to make informed decisions in selecting applications. Subsequently, it navigates and executes actions on their controls, emulating human users to fulfill requests in natural language. The system incorporates a control interaction component, which plays a crucial role in translating actions from GPT-Vision into grounded execution on applications. This feature ensures complete automation without the need for human intervention, thereby establishing it as a comprehensive LAM framework.
@@ 2.2 LLM-based GUI Intelligence
The utilization of multimodal LLM systems Durante et al. (2024) for navigating and controlling GUIs in applications has emerged as a prominent and burgeoning research area. Yan et al., Yan et al. (2023) employed GPT-4V Yang et al. (2023b) to navigate mobile applications by inputting screenshots of GUIs to the LLM, demonstrating state-of-the-art performance across various datasets and human-designed navigation tasks. Similarly, AppAgent Yang et al. (2023a) leverages GPT-4V as smartphone users, enabling them to take actions on mobile applications based on smartphone snapshots, thereby autonomously fulfilling user requests on physical phones. MobileAgent Wang et al. (2024) integrates Optical Character Recognition (OCR) tools to augment the GPT-V employed within a comparable mobile agent designed for task completion on mobile phones. This integration results in a notable enhancement, enabling MobileAgent to achieve completion rates comparable to human performance. On the other hand, CogAgent Hong et al. (2023b) takes a distinct approach by training a dedicated visual language model specialized in GUI understanding and navigation, providing a more tailored solution for various GUI intelligence tasks.
@@ 3.1 UFO in a Nutshell
First, we present the comprehensive architecture of UFO in Figure 2. UFO operates as a dual-agent framework, encompassing (i) an Application Selection Agent (AppAgent) tasked with choosing an application for fulfilling user requests. This agent may also switch to a different application when a request spans multiple applications, and the task is partially completed in the preceding application. Additionally, (ii) an Action Selection Agent (ActAgent) is incorporated, responsible for iteratively executing actions on the selected applications until the task is successfully concluded within a specific application. Both agents leverage the multi-modal capabilities of GPT-Vision to comprehend the application UI and fulfill the user’s request. They utilize a Control Interaction module to ground their actions, enabling tangible impacts on the system.
@@
Upon receiving a user request, the AppAgent undertakes an analysis of the demand. It endeavors to select a fitting application for fulfilling the request from the currently active applications. UFO equips AppAgent with the full desktop screenshot and a list of available applications for reference, facilitating AppAgent’s decision-making process. Subsequently, AppAgent selects an appropriate application and formulates a comprehensive global plan for request completion. This plan is then passed on to the ActAgent.
@@
Once a suitable application is identified, it is brought into focus on the desktop. The ActAgent then initiates actions to fulfill the user request. Before each action selection step, UFO captures screenshots of the current application’s UI window with all available controls annotated. Additionally, UFO records information about each control for ActAgent’s observation. ActAgent is tasked with choosing a control to operate and subsequently selecting a specific action to execute on the chosen control via a control interaction module. This decision is based on ActAgent’s observation, its prior plan, and its operation memory. Following the execution, UFO constructs a local plan for future steps, and proceeds to the next action selection step. This recursive process continues until the user request is successfully completed within the selected application. This concludes one phase of the user request.
@@
In scenarios where the user request spans multiple applications, the ActAgent will delegate the task to the AppAgent for the purpose of switching to a different application once ActAgent completes its tasks on the current one, initiating the second phase of the request. This iterative process continues until all aspects of the user request are fully completed. Users have the option to introduce new requests interactively, prompting UFO to process the new request by repeating the aforementioned process. Upon the successful completion of all user requests, UFO concludes its operation. In the subsequent sections, we delve into the intricate details of each component within the UFO framework.
3.2 AppAgent
Figure 3: An illustration of the AppAgent.
@@ 3.2 AppAgent
The AppAgent bears the responsibility of selecting an active application to fulfill user requests or switching to a new application when necessary. Additionally, AppAgent constructs a comprehensive global plan to orchestrate the entire task, and its architecture is illustrated in Figure 3. AppAgent takes the following information as input:
@@
The provided information, including Desktop Screenshots, Application Information, and Memory, collectively equips AppAgent with a comprehensive set of data to facilitate decision-making. Desktop Screenshots and Application Information enable AppAgent to comprehend the current status and constrain its scope for application selection. On the other hand, Memory serves as a historical record of past request completions, aiding AppAgent in making informed decisions based on prior experiences. This multifaceted input framework enhances AppAgent’s ability to select the most suitable application for fulfilling user requests.
Upon collecting all pertinent information, AppAgent employs GPT-V to generate the following outputs:
@@
Prompting AppAgent to provide its Observation and Thoughts serves a dual purpose. Firstly, it encourages AppAgent to meticulously analyze the current status, offering a detailed explanation for its logic and decision-making process. This not only enhances the logical coherence of its decisions Wei et al. (2022); Ding et al. (2023), but also contributes to improving the overall interpretability of UFO. Secondly, AppAgent determines the status of the task, outputting “FINISH” if it deems the task completed. AppAgent may also leave comments for the user, reporting progress, highlighting potential issues, or addressing any user queries. Once AppAgent identifies the selected application, UFO proceeds to take specific actions within this application to fulfill the user request, with the ActAgent responsible for executing these actions.
3.3 ActAgent
Figure 4: An illustration of the ActAgent.
The ActAgent functions as a downstream entity following AppAgent, responsible for executing specific actions on the selected application to fulfill the user request. Its input and output also exhibit variations compared to AppAgent, as depicted in Figure 4. ActAgent accepts the following input:
  • User Request: The original user query submitted to UFO, identical to the AppAgent.
@@ 3.3 ActAgent
  • Memory: Previous thoughts, comments, actions, and execution results, mirroring the AppAgent.
@@
In contrast to the AppAgent, UFO provides ActAgent with three types of screenshots to aid in its decision-making process. The previous screenshot with the last selected control highlighted in a red rectangle (i.e.,  ) helps comprehend the operation execution in the last step and analyze the impact of the action. The clean screenshot allows for an understanding of the application’s status without annotation obstructions, and the annotated screenshot, labeling each control with a number (e.g., 36), facilitates a better understanding of the function and location of UI elements. Different types of controls are labeled with distinct colors for differentiation.
Moreover, the memory fed into ActAgent serves two key purposes. Firstly, it acts as a reminder for the agent, enabling ActAgent to analyze past actions and reduce the likelihood of repeating actions that proved ineffective. Secondly, it establishes a crucial channel for cross-application communication. Execution results, such as text extracted from a document or a description of an image, are stored in the memory module. ActAgent can selectively incorporate this information for actions that require it, such as composing an email with text from different sources. This augmentation significantly extends the capabilities of UFO.
Given this comprehensive input, ActAgent meticulously analyzes all the information and outputs the following:
@@
  • Comment: Additional comments or information, including a brief progress summary, highlighted points, or changes in the plan, similar to what AppAgent provides.
While some of its output fields may share similarities with AppAgent, UFO determines the next step based on the outputted Status of the task. If the task is not finished, it applies the Function to the selected control, triggering the next state of the application after execution. ActAgent iteratively repeats this process of observing and reacting to the selected application until the user request is fully completed or a switch to a different application is required.
@@ 3.4 Control Interaction
To execute and ground the selected action from ActAgent on the application controls, UFO first detects and translates the action into an executable operation. The Python package pywinauto Bim & Min-shuai (2014) offers valuable tools for inspecting UI controls and performing actions on them. It provides a range of functions and classes for automating the Windows GUI, allowing for programmatic interaction and control of Windows applications. This capability is particularly useful for tasks such as automated testing, scripting, and the automation of repetitive tasks. As the backend, we choose the Windows UI Automation API Dinh et al. (2018) for its robust support in UI inspection and interaction through code.
UFO utilizes pywinauto to inspect all actionable controls of an application, retrieving their precise location and boundary box to facilitate annotations. An example of annotated controls on a PowerPoint GUI is presented in Figure 5, with control information provided by pywinauto. Different colors various control types. Additionally, pywinauto provides rich context for each control, including its name, type, and title, which are crucial information for control and action selection.
@@
Button
A button is a user interface element that users can interact with to trigger an action. Clicking the button typically initiates a specific operation or command.
Edit
An edit control allows users to input and edit text or numeric data. It is commonly used for fields where users can type information, such as textboxes or search bars.
TabItem
A tab item is part of a tab control, organizing content into multiple pages. Users can switch between different tab items to access distinct sets of information or functionalities.
Document
A document control represents a document or a page in a document-view architecture. It is often used to display and manage documents or large blocks of text.
ListItem
A list item is an element within a list control, presenting data in a list format. Users can select and interact with individual items within the list.
MenuItem
A menu item is part of a menu control, providing a list of commands or options. Users can click on menu items to trigger specific actions or navigate through application features.
TreeItem
A tree item is a node within a tree control, organizing information in a hierarchical structure. Users can expand or collapse tree items to navigate through a hierarchical set of data.
ComboBox
A ComboBox is a combination of a text box and a drop-down list. It allows users to either type a value directly into the text box or choose from a predefined list by opening the drop-down menu.
Hyperlink
A Hyperlink enables users to navigate to another location or resource. They are often used to provide easy access to external websites, documents, or specific sections within an application.
ScrollBar
A scroll bar allows users to scroll through content that is larger than the visible area. It provides a way to navigate vertically or horizontally within a window or control.
@@
Click, SetText, GetText, and Scroll are common functions originally supported by pywinauto, covering most daily operations on GUI. Annotate and Summary are customized operations to fulfill special requests for UFO. The former allows re-annotation of the GUI with a more concise list of controls (details in Section 3.5.3), and the latter enables UFO to describe its visual observation in text to fulfill user requests. At every step, ActAgent will choose an action from this available list to execute on the selected UI control in the application. With the control interaction module, UFO evolves into a LAM capable of making tangible impacts on the system.
@@ 3.5.5 Safeguard
Send Action for a Message or Email
Initiating the “Send” action, such as clicking the send button, is considered sensitive as the dispatched message or email becomes irretrievable.
Deleting or Modifying Files and Folders
Operations involving the deletion or modification of files and folders, especially those situated in critical system directories or containing vital user data.
Closing a Window or Application
Closing a window or application is flagged as sensitive due to the potential for data loss or system crashes.
Accessing Webcam or Microphone
Accessing the webcam or microphone without explicit user consent is identified as sensitive to address privacy concerns.
Installing or Uninstalling Software
Actions related to installing or uninstalling software applications are marked as sensitive due to their impact on system configuration and potential security risks.
Browser History or Password Retrieval
Retrieving sensitive user data such as browser history or stored passwords is identified as a sensitive action, posing potential privacy leaks.

### v5

@@ 1 Introduction
The advent of Large Language Models (LLMs) Ouyang et al. (2022); Achiam et al. (2023); Touvron et al. (2023) has showcased revolutionary potential in solving complex problems akin to human reasoning, planning, and collaboration Hao et al. (2023); Ding et al. (2023); Talebirad & Nadiri (2023). This development brings humanity closer to the realization of Artificial General Intelligence (AGI) Liu et al. (2023), offering assistance in various tasks in our daily lives and exhibiting a robust and comprehensive level of ability that was previously considered challenging Chen et al. (2023b); Shao et al. (2023); Jin et al. (2023). The advancement towards more sophisticated multi-modal visual LLMs, exemplified by GPT-Vision Zhang et al. (2023); Yang et al. (2023c), introduces a visual dimension to LLMs, expanding their capabilities to encompass a myriad of visual tasks and broadening their scope to address challenges that require visual capability Wu et al. (2023a); Zheng et al. (2024); Wake et al. (2023).
@@
In this context, the Windows operating system (OS) stands out as representative platform for LAMs, due to its high market share in the daily use of computer systems Adekotujo et al. (2020), the presence of versatile applications and GUIs built upon it Ramler et al. (2018), and the complexity of tasks that necessitate long-term planning and interaction across various applications Stallings (2005). The prospect of having a general intelligent agent that can comprehend user requests in natural language, and autonomously interact with the UIs of applications built on Windows is highly appealing. Despite the evident demand for developing VLM agents tailored for Windows OS to fulfill user requests, this direction remains largely unexplored, as most existing agents predominantly focus on smartphones Yang et al. (2023b) or web applications Zheng et al. (2024). This gap in exploration presents a vast, untapped opportunity to develop a Windows-specific agent.
To bridge this gap, we introduce UFO [Uncaptioned image], a specialized UI-Focused Agent designed for seamless interaction with the Windows OS, leveraging the cutting-edge capabilities of the VLM GPT-Vision Yang et al. (2023c). UFO adopts a dual-agent framework, with each agent analyzing screenshots and extracting information from the GUI to make informed decisions in selecting applications. Subsequently, it navigates and executes actions on their controls, emulating human users to fulfill requests in natural language. The system incorporates a control interaction component, which plays a crucial role in translating actions from GPT-Vision into grounded execution on applications. This feature ensures complete automation without the need for human intervention, thereby establishing it as a comprehensive LAM framework.
@@ 2.2 LLM-based GUI Intelligence
The utilization of multimodal LLM systems Durante et al. (2024) for navigating and controlling GUIs in applications has emerged as a prominent and burgeoning research area. Yan et al., Yan et al. (2023) employed GPT-4V Yang et al. (2023c) to navigate mobile applications by inputting screenshots of GUIs to the LLM, demonstrating state-of-the-art performance across various datasets and human-designed navigation tasks. Similarly, AppAgent yang2023HostAgent leverages GPT-4V as smartphone users, enabling them to take actions on mobile applications based on smartphone snapshots, thereby autonomously fulfilling user requests on physical phones. MobileAgent Wang et al. (2024) integrates Optical Character Recognition (OCR) tools to augment the GPT-V employed within a comparable mobile agent designed for task completion on mobile phones. This integration results in a notable enhancement, enabling MobileAgent to achieve completion rates comparable to human performance. On the other hand, CogAgent Hong et al. (2023b) takes a distinct approach by training a dedicated visual language model specialized in GUI understanding and navigation, providing a more tailored solution for various GUI intelligence tasks.
@@ 3.1 UFO in a Nutshell
First, we present the comprehensive architecture of UFO in Figure 2. UFO operates as a dual-agent framework, encompassing (i) an HostAgent tasked with choosing an application for fulfilling user requests. This agent may also switch to a different application when a request spans multiple applications, and the task is partially completed in the preceding application. Additionally, (ii) an AppAgent is incorporated, responsible for iteratively executing actions on the selected applications until the task is successfully concluded within a specific application. Both agents leverage the multi-modal capabilities of GPT-Vision to comprehend the application UI and fulfill the user’s request. They utilize a Control Interaction module to ground their actions, enabling tangible impacts on the system.
@@
Upon receiving a user request, the HostAgent undertakes an analysis of the demand. It endeavors to select a fitting application for fulfilling the request from the currently active applications. UFO equips HostAgent with the full desktop screenshot and a list of available applications for reference, facilitating HostAgent’s decision-making process. Subsequently, HostAgent selects an appropriate application and formulates a comprehensive global plan for request completion. This plan is then passed on to the AppAgent.
@@
Once a suitable application is identified, it is brought into focus on the desktop. The AppAgent then initiates actions to fulfill the user request. Before each action selection step, UFO captures screenshots of the current application’s UI window with all available controls annotated. Additionally, UFO records information about each control for AppAgent’s observation. AppAgent is tasked with choosing a control to operate and subsequently selecting a specific action to execute on the chosen control via a control interaction module. This decision is based on AppAgent’s observation, its prior plan, and its operation memory. Following the execution, UFO constructs a local plan for future steps, and proceeds to the next action selection step. This recursive process continues until the user request is successfully completed within the selected application. This concludes one phase of the user request.
@@
In scenarios where the user request spans multiple applications, the AppAgent will delegate the task to the HostAgent for the purpose of switching to a different application once AppAgent completes its tasks on the current one, initiating the second phase of the request. This iterative process continues until all aspects of the user request are fully completed. Users have the option to introduce new requests interactively, prompting UFO to process the new request by repeating the aforementioned process. Upon the successful completion of all user requests, UFO concludes its operation. In the subsequent sections, we delve into the intricate details of each component within the UFO framework.
3.2 HostAgent
Figure 3: An illustration of the HostAgent.
@@ 3.2 HostAgent
The HostAgent bears the responsibility of selecting an active application to fulfill user requests or switching to a new application when necessary. Additionally, HostAgent constructs a comprehensive global plan to orchestrate the entire task, and its architecture is illustrated in Figure 3. HostAgent takes the following information as input:
@@
The provided information, including Desktop Screenshots, Application Information, and Memory, collectively equips HostAgent with a comprehensive set of data to facilitate decision-making. Desktop Screenshots and Application Information enable HostAgent to comprehend the current status and constrain its scope for application selection. On the other hand, Memory serves as a historical record of past request completions, aiding HostAgent in making informed decisions based on prior experiences. This multifaceted input framework enhances HostAgent’s ability to select the most suitable application for fulfilling user requests.
Upon collecting all pertinent information, HostAgent employs GPT-V to generate the following outputs:
@@
Prompting HostAgent to provide its Observation and Thoughts serves a dual purpose. Firstly, it encourages HostAgent to meticulously analyze the current status, offering a detailed explanation for its logic and decision-making process. This not only enhances the logical coherence of its decisions Wei et al. (2022); Ding et al. (2023), but also contributes to improving the overall interpretability of UFO. Secondly, HostAgent determines the status of the task, outputting “FINISH” if it deems the task completed. HostAgent may also leave comments for the user, reporting progress, highlighting potential issues, or addressing any user queries. Once HostAgent identifies the selected application, UFO proceeds to take specific actions within this application to fulfill the user request, with the AppAgent responsible for executing these actions.
3.3 AppAgent
Figure 4: An illustration of the AppAgent.
The AppAgent functions as a downstream entity following HostAgent, responsible for executing specific actions on the selected application to fulfill the user request. Its input and output also exhibit variations compared to HostAgent, as depicted in Figure 4. AppAgent accepts the following input:
  • User Request: The original user query submitted to UFO, identical to the HostAgent.
@@ 3.3 AppAgent
  • Memory: Previous thoughts, comments, actions, and execution results, mirroring the HostAgent.
@@
In contrast to the HostAgent, UFO provides AppAgent with three types of screenshots to aid in its decision-making process. The previous screenshot with the last selected control highlighted in a red rectangle (i.e.,  ) helps comprehend the operation execution in the last step and analyze the impact of the action. The clean screenshot allows for an understanding of the application’s status without annotation obstructions, and the annotated screenshot, labeling each control with a number (e.g., 36) using Set-of-Mark (SoM) Yang et al. (2023a), facilitates a better understanding of the function and location of UI elements. Different types of controls are labeled with distinct colors for differentiation.
Moreover, the memory fed into AppAgent serves two key purposes. Firstly, it acts as a reminder for the agent, enabling AppAgent to analyze past actions and reduce the likelihood of repeating actions that proved ineffective. Secondly, it establishes a crucial channel for cross-application communication. Execution results, such as text extracted from a document or a description of an image, are stored in the memory module. AppAgent can selectively incorporate this information for actions that require it, such as composing an email with text from different sources. This augmentation significantly extends the capabilities of UFO.
Given this comprehensive input, AppAgent meticulously analyzes all the information and outputs the following:
@@
  • Comment: Additional comments or information, including a brief progress summary, highlighted points, or changes in the plan, similar to what HostAgent provides.
While some of its output fields may share similarities with HostAgent, UFO determines the next step based on the outputted Status of the task. If the task is not finished, it applies the Function to the selected control, triggering the next state of the application after execution. AppAgent iteratively repeats this process of observing and reacting to the selected application until the user request is fully completed or a switch to a different application is required.
@@ 3.4 Control Interaction
To execute and ground the selected action from AppAgent on the application controls, UFO first detects and translates the action into an executable operation. The Python package pywinauto Bim & Min-shuai (2014) offers valuable tools for inspecting UI controls and performing actions on them. It provides a range of functions and classes for automating the Windows GUI, allowing for programmatic interaction and control of Windows applications. This capability is particularly useful for tasks such as automated testing, scripting, and the automation of repetitive tasks. As the backend, we choose the Windows UI Automation API Dinh et al. (2018) for its robust support in UI inspection and interaction through code.
UFO utilizes pywinauto to inspect all actionable controls of an application, retrieving their precise location and boundary box to facilitate annotations with SoM Yang et al. (2023a). An example of annotated controls on a PowerPoint GUI is presented in Figure 5, with control information provided by pywinauto. Different colors various control types. Additionally, pywinauto provides rich context for each control, including its name, type, and title, which are crucial information for control and action selection.
@@
Button A button is a user interface element that users can interact with to trigger an action. Clicking the button typically initiates a specific operation or command.
Edit An edit control allows users to input and edit text or numeric data. It is commonly used for fields where users can type information, such as textboxes or search bars.
TabItem A tab item is part of a tab control, organizing content into multiple pages. Users can switch between different tab items to access distinct sets of information or functionalities.
Document A document control represents a document or a page in a document-view architecture. It is often used to display and manage documents or large blocks of text.
ListItem A list item is an element within a list control, presenting data in a list format. Users can select and interact with individual items within the list.
MenuItem A menu item is part of a menu control, providing a list of commands or options. Users can click on menu items to trigger specific actions or navigate through application features.
TreeItem A tree item is a node within a tree control, organizing information in a hierarchical structure. Users can expand or collapse tree items to navigate through a hierarchical set of data.
ComboBox A ComboBox is a combination of a text box and a drop-down list. It allows users to either type a value directly into the text box or choose from a predefined list by opening the drop-down menu.
Hyperlink A Hyperlink enables users to navigate to another location or resource. They are often used to provide easy access to external websites, documents, or specific sections within an application.
ScrollBar A scroll bar allows users to scroll through content that is larger than the visible area. It provides a way to navigate vertically or horizontally within a window or control.
@@
Click, SetText, GetText, and Scroll are common functions originally supported by pywinauto, covering most daily operations on GUI. Annotate and Summary are customized operations to fulfill special requests for UFO. The former allows re-annotation of the GUI with a more concise list of controls (details in Section 3.5.3), and the latter enables UFO to describe its visual observation in text to fulfill user requests. At every step, AppAgent will choose an action from this available list to execute on the selected UI control in the application. With the control interaction module, UFO evolves into a LAM capable of making tangible impacts on the system.
@@ 3.5.5 Safeguard
Send Action for a Message or Email Initiating the “Send” action, such as clicking the send button, is considered sensitive as the dispatched message or email becomes irretrievable.
Deleting or Modifying Files and Folders Operations involving the deletion or modification of files and folders, especially those situated in critical system directories or containing vital user data.
Closing a Window or Application Closing a window or application is flagged as sensitive due to the potential for data loss or system crashes.
Accessing Webcam or Microphone Accessing the webcam or microphone without explicit user consent is identified as sensitive to address privacy concerns.
Installing or Uninstalling Software Actions related to installing or uninstalling software applications are marked as sensitive due to their impact on system configuration and potential security risks.
Browser History or Password Retrieval Retrieving sensitive user data such as browser history or stored passwords is identified as a sensitive action, posing potential privacy leaks.